OpenStudy (anonymous):

FRACTIONS NEED ANSWERS: 1: 4 5/6 + 7 7/24 2: 8 1/4 - 3 1/3 3: 5 3/7 + 4 1/2 4: 3 5/12 + 7 2/15

OpenStudy (anonymous):

YES THESE ARE MIXED NUMBERS

OpenStudy (anonymous):

i'll do the first one for you: 1 4+ 7 = 11 find lcm of 6 and 24 = 24: 5/6 + 7/24 = 20/24 + 7/24 = 27/24 = = 1 3/24 = 1 1/8 11 + 1 1/8 = 12 1/8